Soft drink cans are filled by an automated filling machine and the standard deviation is 0.5 fluid ounce. Assume that the fill volumes of the cans are independent, normal random variables.

Round your answers to four decimal places (e.g. 98.7654).

If the mean fill volume is 12.1 fluid ounces, what should the standard deviation of fill volume equal so that the probability that the average of 100 cans is less than 12 fluid ounces is 0.005?
Answer #1

Answer Given that, mean, x 12.1 fluid ounces n 100 P(x < 12) = 0.005 z. score :-2.576 z = x-x 12-12.1 -2.576 = 10 - 2.576 -0.